# Electromagnetic Property Sensing in ISAC with Multiple Base Stations: Algorithm, Pilot Design, and Performance Analysis

### v1

@@ Abstract
Integrated sensing and communication (ISAC) has opened up numerous game-changing opportunities for future wireless systems. In this paper, we develop a novel scheme that utilizes orthogonal frequency division multiplexing (OFDM) pilot signals to sense the electromagnetic (EM) property of the target and thus identify the materials of the target. Specifically, we first establish an EM wave propagation model with Maxwell equations, where the EM property of the target is captured by a closed-form expression of the channel. We then build the mathematical model for the relative permittivity and conductivity distribution (RPCD) within a predetermined region of interest shared by multiple base stations (BSs). Based on the EM wave propagation model, we propose an EM property sensing method, in which the RPCD can be reconstructed from compressive sensing techniques that exploits the joint sparsity structure of the EM property vector. We then develop a fusion algorithm to combine data from multiple BSs, which can enhance the reconstruction accuracy of EM property by efficiently integrating diverse measurements. Moreover, the fusion is performed at the feature level of RPCD and features low transmission overhead. We further design the pilot signals that can minimize the mutual coherence of the equivalent channels and enhance the diversity of incident EM wave patterns. Simulation results demonstrate the efficacy of the proposed method in achieving high-quality RPCD reconstruction and accurate material classification.
@@ I Introduction
Recently, the emergence of integrated sensing and communication (ISAC) has unlocked a multitude of revolutionary possibilities for the forthcoming wireless sensing systems. ISAC facilitates the sharing of the limited radio spectrum between sensing and communications (S&C) systems and thereby significantly reduces resource costs [1, 2, 3, 4]. In contrast to the conventional approach of separate functionalities for S&C, ISAC methodology brings two distinct advantages. Firstly, by enabling the joint utilization of constrained resources such as spectrum, energy, and hardware platforms, ISAC enhances efficiency for both S&C and thus yields integration benefits. Secondly, the collaborative synergy between S&C fosters coordination gains and thereby amplifies the overall performance. Consequently, ISAC has the potential to improve the capabilities of both S&C functionalities in the foreseeable future [5, 6, 7, 8]. Thanks to its myriad benefits, ISAC would play a pivotal role in various forthcoming applications, including intelligent connected vehicles, internet of things (IoT), and smart homes and cities [9, 10, 11, 12].
ISAC with multiple base stations (BSs) has emerged as a pivotal area of research, which focuses on the coordinated use of resources across a network of BSs to simultaneously enhance S&C capabilities [13, 14, 15]. This approach is underpinned by the principle of leveraging shared information and coordinated actions among multiple BSs. Recent advancements have explored various facets of ISAC with multiple BSs, including power control [16], cooperative sensing [17], and beamforming strategies [18], which collectively optimize the BSs’ ability to sense the environment in a more accurate way while maintaining robust communication links. These efforts are crucial to realize the next generation of wireless networks, where the seamless integration of S&C will lead to innovative applications in areas such as autonomous vehicles, smart cities, and advanced surveillance systems [19, 20, 21, 22].
@@
Traditional methods sense EM property via infrared detection, which utilizes the unique infrared emissivity property of each material [25]. However, these methods typically require expensive and specialized equipment that is difficult to produce and sustain [26]. Additionally, infrared detection is hindered by its limited ability to penetrate targets, especially when the surface is covered with camouflage, which makes the conventional methods ineffective. On the other hand, the ability of relatively low-frequency EM waves to penetrate various objects makes EM wave detection in communication frequency band a practical and promising alternative for reliable EM property sensing [27]. Since the EM property of target materials is implicitly embedded within the channel state information (CSI) from transmitters to receivers, it is possible to sense EM property via ISAC technologies.
Although ISAC has demonstrated significant achievements in localization [28], tracking [29], imaging [30], as well as various other applications [5, 7, 31, 9], EM property sensing within the ISAC framework has been barely studied. To the best of the authors’ knowledge, only one work [32] implements EM property sensing under single-BS ISAC systems. However, the estimation accuracy and the resolution under a single-BS ISAC system are not satisfactory.
@@
    We propose a fusion algorithm to combine data from multiple BSs at the RPCD feature level, which enhances the reconstruction accuracy of EM property by efficiently integrating diverse measurements and boasts small information redundancy as well as low transmission overhead.
@@
Notations: Boldface denotes a vector or a matrix; $j$ corresponds to the imaginary unit; $(\cdot)^{H}$, $(\cdot)^{T}$, and $(\cdot)^{*}$ represent Hermitian, transpose, and conjugate, respectively; $\circ$ denotes the Khatri-Rao product; $\odot$ denotes the Hadamard product; $\otimes$ denotes the Kronecker product; $\mathrm{vec}(\cdot)$ denotes the vectorization operation; $\textrm{diag}(\mathbf{a})$ denotes the diagonal matrix whose diagonal elements are the elements of $\mathbf{a}$; $\mathbf{I}$, $\mathbf{1}$, and $\mathbf{0}$ denote the identity matrix, the all-ones vector, and the all-zeros vector with compatible dimensions; $\left\|\mathbf{a}\right\|_{2}$ denotes $\ell 2$-norm of the vector $\mathbf{a}$; $\left|\mathbf{a}\right|$ denotes the element-wise absolute value of the complex vector $\mathbf{a}$; $\Re(\cdot)$ and $\Im(\cdot)$ denote the real and imaginary part of complex vectors or matrices, respectively; $\text{Tr}(\cdot)$ denotes the trace of a matrix; $\textbf{prox}_{f(\cdot)}$ denotes the proximal mapping of the function $f(\cdot)$; $\left\|\mathbf{A}\right\|_{F}$ denotes Frobenius-norm of the matrix $\mathbf{A}$; $\mathbf{A}\succeq\mathbf{0}$ indicates that the matrix $\mathbf{A}$ is positive semi-definite; the distribution of a circularly symmetric complex Gaussian random vector with zero mean and covariance matrix $\mathbf{A}$ is denoted as $\mathcal{CN}(\mathbf{0},\mathbf{A})$.
@@ II-A Signal Model
Consider an uplink OFDM communication scenario with $U$ mobile UEs and $L$ BSs. The $L$ BSs collaboratively provide EM property sensing service for a single target and are connected using cables by a central processing unit. Assume that each UE is equipped with $N_{t}$ transmitting antennas, and each BS is equipped with $N_{r}$ receiving antennas. The scenario is also known as the multi-static ISAC system. The UEs all adopt fully digital precoding structure where the number of RF chains $N_{RF}$ is equal to the number of antennas $N_{t}$. To facilitate EM property sensing, we utilize pilot transmissions over $K$ subcarriers. The central frequency of the OFDM signals is represented by $f_{c}$, and the interval between subcarriers is denoted by $\Delta f$.
In order to avoid interference between multiple UEs, we assume the ISAC system utilizes the time-division duplex (TDD) mode, i.e., only one UE sends the pilot signals at one time. Additionally, it is assumed that each UE transmits $I$ pilot symbols at each subcarrier.
@@
Suppose the target scatters the pilot signals from the UE to the BSs as shown in Fig. 1. Since only the signals scattered by the target carry the information of its EM property, the UE may purposely send the pilot signals towards the target by beamforming. The overall ISAC channel from the $u$-th UE to the $l$-th BS at the $k$-th subcarrier thus only consists of the target scattering path channel $\hat{\mathbf{H}}_{k,l,u}\in\mathbb{C}^{N_{r}\times N_{t}}$. Thus, the signals received by the $l$-th BS can be formulated as:
$\tilde{\mathbf{y}}_{k,l,u}=\hat{\mathbf{H}}_{k,l,u}\mathbf{w}_{k,u}+\tilde{\mathbf{n}}_{k,l},$ (1)
where $\mathbf{w}_{k,u}\in\mathbb{C}^{N_{t}\times 1}$ is the digital pilot symbols on the $k$-th subcarrier, and $\tilde{\mathbf{n}}_{k,l}\sim\mathcal{CN}\left(\mathbf{0}_{N_{r}},\sigma_{k,l}^{2}\mathbf{I}_{N_{r}}\right)$ is the complex Gaussian noise at the $l$-th BS on the $k$-th subcarrier. We will demonstrate that $\mathbf{w}_{k,u}$ should be designed in a varying way across the subcarriers and hence a fully digital system is required.
Assume prior knowledge has determined that the target is contained within a certain region to be sensed, denoted by $D$, which can be discretized into $M$ sampling points. Let the vectors $\mathbf{E}_{k}^{i,D}\in\mathbb{C}^{M\times 1}$ and $\mathbf{J}_{k}^{D}\in\mathbb{C}^{M\times 1}$ collect the incident electric field and the equivalent contrast source at all $M$ sampling points in $D$, respectively [34, 35, 36]. Define $\mathbf{H}_{1,k,u}\in\mathbb{C}^{M\times N_{t}}$ as the channel matrix from the $u$-th UE to the target that maps $\mathbf{w}_{k,u}$ to $\mathbf{E}_{k}^{i,D}$, and define $\mathbf{H}_{2,k,l}\in\mathbb{C}^{N_{r}\times M}$ as the channel matrix from the target to the $l$-th BS path that maps $\mathbf{J}_{k}^{D}$ to $\tilde{\mathbf{y}}_{k,l}$. The ISAC channel from the UE through region $D$ to the $l$-th BS can then be described as:
@@ II-B EM Wave Propagation Model
Let the vectors $\mathbf{E}_{k}^{s,D}\in\mathbb{C}^{M\times 1}$, $\mathbf{E}_{k}^{t,D}\in\mathbb{C}^{M\times 1}$, and $\boldsymbol{\chi}_{k}\in\mathbb{C}^{M\times 1}$ collect the scattering electric field, the total electric field, and $\chi_{k}(\mathbf{r})$ on all $M$ sampling points in $D$, respectively. Let $\mathbf{G}_{k}\in\mathbb{C}^{M\times M}$ represent the matrix obtained by discretizing the integral kernel $k_{k}^{2}G_{k}(\mathbf{r},\mathbf{r}^{\prime})$ in (4) using the method of moments (MoM). In $\mathbf{G}_{k}$, both $\mathbf{r}$ and $\mathbf{r}^{\prime}$ correspond to the discretized sampling points within region $D$.
According to [32], the closed-form expression of $\mathbf{X}_{k}$ in (2) can be written as:
@@
Denote $\tilde{\mathbf{W}}_{k,u}=\left[\mathbf{w}_{k,u,1},\mathbf{w}_{k,u,2},\cdots,\mathbf{w}_{k,u,I}\right]\in\mathbb{C}^{N_{t}\times I}$ as the digital transmitter beamforming matrix stacked by time at the $k$-th subcarrier. Then the overall received pilot signals $\mathbf{Y}_{k,l,u}\in\mathbb{C}^{N_{r}\times I}$ can be formulated in a compact form as:
@@
$=\mathbf{H}_{2,k,l}\textrm{diag}\left({\boldsymbol{\chi}_{k}}\right)[\mathbf{I}-\mathbf{G}_{k}\textrm{diag}\left(\boldsymbol{\chi}_{k}\right)]^{-1}\mathbf{H}_{1,k,u}\tilde{\mathbf{W}}_{k,u}$
@@
=$\overset{(a)}{=}\!\!\left[\left(\bar{\mathbf{H}}_{1,k}^{\top}\left[\mathbf{I}-\mathbf{G}_{k}\textrm{diag}\left({\boldsymbol{\chi}_{k}}\right)\right]^{-\top}\right)\circ\mathbf{H}_{2,k,l}\right]\boldsymbol{\chi}_{k}+\mathbf{n}_{k,l},$ (7)
@@
$\mathbf{D}_{k,l}$ =$\overset{\Delta}{=}\left\{\bar{\mathbf{H}}_{1,k}^{\top}\left[\mathbf{I}-\mathbf{G}_{k}\textrm{diag}\left({\boldsymbol{\chi}_{k}}\right)\right]^{-\top}\right\}\circ\mathbf{H}_{2,k,l}.$ (8)

### v2

@@ Abstract
Integrated sensing and communication (ISAC) has opened up numerous game-changing opportunities for future wireless systems. In this paper, we develop a novel scheme that utilizes orthogonal frequency division multiplexing (OFDM) pilot signals to sense the electromagnetic (EM) property of the target and thus identify the materials of the target. Specifically, we first establish an EM wave propagation model with Maxwell equations, where the EM property of the target is captured by a closed-form expression of the channel. We then build the mathematical model for the relative permittivity and conductivity distribution (RPCD) within a predetermined region of interest shared by multiple base stations (BSs). By leveraging the Lippmann-Schwinger equation, we propose an EM property sensing method that reconstructs the RPCD using compressive sensing techniques. This approach exploits the joint sparsity of the EM property vector, which enables the proposed method to effectively handle the high dimensionality and ill-posed nature of the inverse scattering problem. We then develop a fusion algorithm to combine data from multiple BSs, which can enhance the reconstruction accuracy of EM property by efficiently integrating diverse measurements. Moreover, the fusion is performed at the feature level of RPCD and features low transmission overhead. We further design the pilot signals that can minimize the mutual coherence of the equivalent channels and enhance the diversity of incident EM wave patterns. Simulation results demonstrate the efficacy of the proposed method in achieving high-quality RPCD reconstruction and accurate material classification.
@@ I Introduction
Recently, the emergence of integrated sensing and communication (ISAC) has unlocked a multitude of revolutionary possibilities for the forthcoming wireless sensing systems. ISAC facilitates the sharing of the limited radio spectrum between sensing and communications (S&C) systems and thereby significantly reduces resource costs [1, 2, 3]. In contrast to the conventional approach of separate functionalities for S&C, ISAC methodology brings two distinct advantages. Firstly, by enabling the joint utilization of constrained resources such as spectrum, energy, and hardware platforms, ISAC enhances efficiency for both S&C and thus yields integration benefits. Secondly, the collaborative synergy between S&C fosters coordination gains and thereby amplifies the overall performance. Consequently, ISAC has the potential to improve the capabilities of both S&C functionalities in the foreseeable future. Thanks to its myriad benefits, ISAC would play a pivotal role in various forthcoming applications, including intelligent connected vehicles, internet of things (IoT), and smart homes and cities [4, 5, 6].
ISAC with multiple base stations (BSs) has emerged as a crucial area of research, which focuses on the coordinated use of resources across a network of BSs to simultaneously enhance S&C capabilities [7, 8, 9]. This approach leverages shared information and coordinated actions among BSs, optimizing performance through advancements in symbol-level cooperative sensing [10], joint fronthaul compression and beamforming design [11], and UAV trajectory design for coordinated beamforming [12]. Shared information among BSs collectively improves environmental sensing accuracy at the cost of large transmission overhead between BSs and the central processing unit. The cooperation of multiple BSs is vital for the next-generation networks, enabling applications like autonomous vehicles, smart cities, and advanced surveillance systems, where ISAC technologies offer both high-resolution sensing and low-latency communications [13, 14, 15, 16].
@@
Traditional methods sense EM property via infrared detection, which utilizes the unique infrared emissivity property of each material [19]. However, these methods typically require expensive and specialized equipment that is difficult to produce and sustain [20]. Additionally, infrared detection is hindered by its limited ability to penetrate targets, especially when the surface is covered with camouflage, which makes the conventional methods ineffective. The low-frequency EM wave is a promising alternative for reliable EM property sensing because they can penetrate various objects more effectively than infrared waves [21]. Due to their longer wavelengths, low-frequency EM waves interact less with small particles and surface irregularities, reducing absorption and scattering. Unlike infrared detection, low-frequency EM waves in communication band enable practical sensing by leveraging the channel state information (CSI) embedded in transmitted signals, which allows ISAC technologies to identify EM property reliably without the need for specialized equipment.
Although ISAC has demonstrated significant achievements in localization [22], tracking [23], imaging [24], as well as various other applications [1, 3, 25], EM property sensing within the ISAC framework has been barely studied. To the best of the authors’ knowledge, only one work [26] implements EM property sensing under single-BS ISAC systems. However, the estimation accuracy and the resolution under a single-BS ISAC system are not satisfactory, and a single BS may not cover a large sensing region. Although multi-BS ISAC systems offer significant potential to improve sensing accuracy by combining observations from different locations, they face challenges in efficiently fusing distributed data with limited transmission overhead. Additionally, [26] only discusses the reduction of pilot mutual coherence within the same subcarrier. Reducing pilot mutual coherence across subcarriers remains unexplored and could enhance sensing reliability and spectrum efficiency in large-scale systems.
@@
    We propose a fusion algorithm to combine data from multiple BSs at the RPCD feature level, which reduces the error of RPCD reconstruction by efficiently integrating signals on different subcarriers and boasts relatively low transmission overhead.
@@
Notations: Boldface denotes a vector or a matrix; $j$ corresponds to the imaginary unit; $(\cdot)^{H}$, $(\cdot)^{T}$, and $(\cdot)^{*}$ represent Hermitian, transpose, and conjugate, respectively; $\circ$ denotes the Khatri-Rao product; $\odot$ denotes the Hadamard product; $\otimes$ denotes the Kronecker product; $\mathrm{vec}(\cdot)$ denotes the vectorization operation; $\textrm{diag}(\mathbf{a})$ denotes the diagonal matrix whose diagonal elements are the elements of $\mathbf{a}$; $\mathbf{I}$, $\mathbf{1}$, and $\mathbf{0}$ denote the identity matrix, the all-ones vector, and the all-zeros vector with compatible dimensions; $\left\|\mathbf{a}\right\|_{2}$ denotes $\ell 2$-norm of the vector $\mathbf{a}$; $\left|\mathbf{a}\right|$ denotes the element-wise absolute value of the complex vector $\mathbf{a}$; $\Re(\cdot)$ and $\Im(\cdot)$ denote the real and imaginary part of complex vectors or matrices, respectively; $\text{Tr}(\cdot)$ denotes the trace of a matrix; $\textbf{prox}_{f(\cdot)}$ denotes the proximal mapping of the function $f(\cdot)$; $\left\|\mathbf{A}\right\|_{F}$ denotes Frobenius-norm of the matrix $\mathbf{A}$; $\mathbf{A}\succeq\mathbf{0}$ indicates that the matrix $\mathbf{A}$ is positive semi-definite; the distribution of a circularly symmetric complex Gaussian (CSCG) random vector with zero mean and covariance matrix $\mathbf{A}$ is denoted as $\mathcal{CN}(\mathbf{0},\mathbf{A})$.
@@ II-A Signal Model
Consider an uplink OFDM communication scenario with $U$ mobile UEs and $L$ BSs. The $L$ BSs collaboratively provide EM property sensing service for a single target and are connected using cables by a central processing unit (CPU). Assume that each UE is equipped with $N_{t}$ transmitting antennas, and each BS is equipped with $N_{r}$ receiving antennas. The scenario is also known as the multi-static ISAC system. The UEs all adopt fully digital precoding structure where the number of RF chains $N_{RF}$ is equal to the number of antennas $N_{t}$. To facilitate EM property sensing, we utilize pilot transmissions over $K$ subcarriers. The central frequency of the OFDM signals is represented by $f_{c}$, and the interval between subcarriers is denoted by $\Delta f$.
In order to avoid interference between multiple UEs, we assume the ISAC system utilizes the time-division duplex (TDD) mode, i.e., only one UE sends the pilot signals at one time. Additionally, it is assumed that each UE transmits $I$ pilot symbols on each subcarrier.
@@
Suppose the target scatters the pilot signals from the UE to the BSs as shown in Fig. 1. Denote the target scattering path channel from the $u$-th UE to the $l$-th BS on the $k$-th subcarrier as $\hat{\mathbf{H}}_{k,l,u}\in\mathbb{C}^{N_{r}\times N_{t}}$ and the direct path channel from the $u$-th UE to the $l$-th BS on the $k$-th subcarrier as $\hat{\mathbf{H}}_{k,l,u}^{d}\in\mathbb{C}^{N_{r}\times N_{t}}$. Then, the signals received by the $l$-th BS can be formulated as
$\hat{\mathbf{y}}_{k,l,u}=\hat{\mathbf{H}}_{k,l,u}\mathbf{w}_{k,u}x_{k,u}+\hat{\mathbf{H}}_{k,l,u}^{d}\mathbf{w}_{k,u}x_{k,u}+\hat{\mathbf{n}}_{k,l},$ (1)
where $\mathbf{w}_{k,u}\in\mathbb{C}^{N_{t}\times 1}$ is the digital beamforming vector on the $k$-th subcarrier, and $\hat{\mathbf{n}}_{k,l}\sim\mathcal{CN}\left(\mathbf{0}_{N_{r}},\sigma_{k,l}^{2}\mathbf{I}_{N_{r}}\right)$ is the complex Gaussian noise at the $l$-th BS on the $k$-th subcarrier. The pilot symbol $x_{k,u}$ can be set as $1$ without loss of generality. Since only the signals scattered by the target carry the information of its EM property, the BS may purposely extract the received signals arriving from the target’s direction by beamforming. Assume prior knowledge has determined that the target is contained within a certain region to be sensed, represented by $D$. Denote $[\theta_{l}^{min},\theta_{l}^{max}]$ as the angular spread of region $D$ when observed from the $l$-th BS. Define the steering vector $\mathbf{a}_{k}(\theta_{l})\in\mathbb{C}^{N_{r}\times 1}$ for $\theta_{l}$ as $\mathbf{a}(\theta_{l})=\frac{1}{\sqrt{N_{r}}}\!\!\begin{bmatrix}1,e^{-j\frac{2\pi}{\lambda_{k}}d_{l}\sin(\theta_{l})},e^{-j\frac{2\pi}{\lambda_{k}}2d_{l}\sin(\theta_{l})},\!\!\cdots\!\!,e^{-j\frac{2\pi}{\lambda_{k}}(N_{r}-1)d_{l}\sin(\theta_{l})}\!\!\!\end{bmatrix}^{\top}\!\!\!\!,$ where $d_{l}$ is the inter-antenna spacing. To extract the signals from the target scattering path, we apply receiver beamforming at the BS with the following beamforming matrix:
$\mathbf{P}_{k,l}=\frac{\int_{\theta_{l}^{min}}^{\theta_{l}^{max}}\mathbf{a}_{k}(\theta_{l})\mathbf{a}_{k}(\theta_{l})^{H}\mathrm{d}\theta_{l}}{\theta_{l}^{max}-{\theta_{l}^{min}}}.$ (2)
If the $u$-th UE, the target, and the $l$-th BS are not on the same line, then there is $\mathbf{P}_{k,l}\hat{\mathbf{H}}_{k,l,u}^{d}\mathbf{w}_{k,u}\approx\mathbf{0}$, especially when the number of the antennas at the BS is large as is true in the coming 6G era. Thus, the received signals after beamforming effectively isolate the direct path of the $u$-th UE and can be simplified to
$\tilde{\mathbf{y}}_{k,l,u}\overset{\Delta}{=}\mathbf{P}_{k,l}\hat{\mathbf{y}}_{k,l,u}=\mathbf{P}_{k,l}\hat{\mathbf{H}}_{k,l,u}\mathbf{w}_{k,u}+\tilde{\mathbf{n}}_{k,l},$ (3)
where $\tilde{\mathbf{n}}_{k,l}\overset{\Delta}{=}\mathbf{P}_{k,l}\hat{\mathbf{n}}_{k,l}$. If the $u$-th UE, the target, and the $l$-th BS happen to be on the same line, then the other UEs can be used to sense the target.
Remark 1.
We will demonstrate that $\mathbf{w}_{k,u}$ should be designed in a varying way across the subcarriers and hence a fully digital system is required.
Suppose we uniformly discretize region $D$ into $M$ sampling points. Let the vectors $\mathbf{E}_{k}^{i,D}\in\mathbb{C}^{M\times 1}$ and $\mathbf{J}_{k}^{D}\in\mathbb{C}^{M\times 1}$ collect the incident electric field and the equivalent contrast source at all $M$ sampling points within $D$, respectively [30, 31, 32]. Define $\mathbf{H}_{1,k,u}\in\mathbb{C}^{M\times N_{t}}$ as the channel matrix from the $u$-th UE to the target that maps $\mathbf{w}_{k,u}$ to $\mathbf{E}_{k}^{i,D}$, and define $\mathbf{H}_{2,k,l}\in\mathbb{C}^{N_{r}\times M}$ as the channel matrix from the target to the $l$-th BS path that maps $\mathbf{J}_{k}^{D}$ to $\tilde{\mathbf{y}}_{k,l}$. The ISAC channel from the UE through region $D$ to the $l$-th BS can then be described as:
@@
The matrices $\mathbf{H}_{1,k,u}$ and $\mathbf{H}_{2,k,l}$ are also known as the radiation operator matrices and are solely related to the physical laws of EM propagation [33, 30]. Since we assume that the positions of the receiver, the region of interest $D$, and the transmitter are known, $\mathbf{H}_{1,k,u}$ and $\mathbf{H}_{2,k,l}$ can be calculated by the computational EM methods such as method of moments (MoM) [30], [34], finite element method (FEM) [35], and finite-difference time-domain (FDTD) [36]. In this paper, we use MoM to compute the radiation operator matrices, which is also widely used in ground penetrating radar [37, 38, 39].
@@ II-B EM Wave Propagation Model
Let the vectors $\mathbf{E}_{k}^{s,D}\in\mathbb{C}^{M\times 1}$, $\mathbf{E}_{k}^{t,D}\in\mathbb{C}^{M\times 1}$, and $\boldsymbol{\chi}_{k}\in\mathbb{C}^{M\times 1}$ collect the scattering electric field, the total electric field, and $\chi_{k}(\mathbf{r})$ on all $M$ sampling points within $D$, respectively. Let $\mathbf{G}_{k}\in\mathbb{C}^{M\times M}$ represent the matrix obtained by discretizing the integral kernel $k_{k}^{2}G_{k}(\mathbf{r},\mathbf{r}^{\prime})$ in (6) using the method of moments (MoM). In $\mathbf{G}_{k}$, both $\mathbf{r}$ and $\mathbf{r}^{\prime}$ correspond to the discretized sampling points within region $D$.
The total electric field within $D$ can be obtained by adding the incident field and the scattering field, which means $\mathbf{E}_{k}^{t,D}$ can be written in the discretized form of (6) specialized with $\mathbf{r}\in D$ as:
$\mathbf{E}_{k}^{t,D}=\mathbf{E}_{k}^{i,D}+\mathbf{E}_{k}^{s,D}=\mathbf{E}_{k}^{i,D}+\mathbf{G}_{k}\textrm{diag}\left({\boldsymbol{\chi}_{k}}\right)\mathbf{E}_{k}^{t,D}.$ (7)
In order to yield a closed-form expression of $\mathbf{E}_{k}^{t,D}$, we reformulate (7) in the nonlinear form with respect to $\textrm{diag}\left({\boldsymbol{\chi}_{k}}\right)$ and then obtain $\mathbf{J}_{k}^{D}$ as:
$\mathbf{E}_{k}^{t,D}=(\mathbf{I}-\mathbf{G}_{k}\textrm{diag}\left({\boldsymbol{\chi}_{k}}\right))^{-1}\mathbf{E}_{k}^{i,D},$ (8)
$\mathbf{J}_{k}^{D}=\textrm{diag}\left({\boldsymbol{\chi}_{k}}\right)\mathbf{E}_{k}^{t,D}=\textrm{diag}\left({\boldsymbol{\chi}_{k}}\right)(\mathbf{I}-\mathbf{G}_{k}\textrm{diag}\left({\boldsymbol{\chi}_{k}}\right))^{-1}\mathbf{E}_{k}^{i,D}.$ (9)
According to (4) and (9), the closed-form expression of $\mathbf{X}_{k}$ in (4) can be written as:
@@
Let $\mathbf{w}_{k,u,i},i=1,\cdots,I$, denote $\mathbf{w}_{k,u}$ of (1) in the $i$-th time slot. Denote $\tilde{\mathbf{W}}_{k,u}=\left[\mathbf{w}_{k,u,1},\mathbf{w}_{k,u,2},\cdots,\mathbf{w}_{k,u,I}\right]\in\mathbb{C}^{N_{t}\times I}$ as the digital transmitter beamforming matrix stacked by time on the $k$-th subcarrier. Then the overall received pilot signals $\mathbf{Y}_{k,l,u}\in\mathbb{C}^{N_{r}\times I}$ can be formulated in a compact form as:
@@
$=\mathbf{P}_{k,l}\mathbf{H}_{2,k,l}\textrm{diag}\left({\boldsymbol{\chi}_{k}}\right)[\mathbf{I}-\mathbf{G}_{k}\textrm{diag}\left(\boldsymbol{\chi}_{k}\right)]^{-1}\mathbf{H}_{1,k,u}\tilde{\mathbf{W}}_{k,u}$
@@
=$\overset{(a)}{=}\!\!\left[\left(\bar{\mathbf{H}}_{1,k}^{\top}\left[\mathbf{I}-\mathbf{G}_{k}\textrm{diag}\left({\boldsymbol{\chi}_{k}}\right)\right]^{-\top}\right)\circ(\mathbf{P}_{k,l}\mathbf{H}_{2,k,l})\right]\boldsymbol{\chi}_{k}+\mathbf{n}_{k,l},$ (12)
@@
$\mathbf{D}_{k,l}$ =$\overset{\Delta}{=}\left\{\bar{\mathbf{H}}_{1,k}^{\top}\left[\mathbf{I}-\mathbf{G}_{k}\textrm{diag}\left({\boldsymbol{\chi}_{k}}\right)\right]^{-\top}\right\}\circ(\mathbf{P}_{k,l}\mathbf{H}_{2,k,l}).$ (13)
@@
Remark 2.
Note that the existing multistatic radar system differs significantly from the system presented in this paper. In radar systems, signals are both transmitted and received by radars, whereas in this paper, mobile UEs transmit the signals and the fixed BSs receive the signals. UEs and BSs vary a lot in their numbers, sizes of antenna arrays, and distances to the target. A lot of UEs surrounding the target can provide more views than only using the BSs. The transmit pilot design is also tailored for UEs with relatively small size of antenna arrays and with relatively close distance to the target. Additionally, traditional multistatic radar imaging systems focus solely on target shape, without considering the distribution of dielectric constants and conductivity. The system most similar to that in this paper is the multi-antenna EM imaging technology [30, 31, 42]. However, in the multi-antenna EM imaging technology, densely arranged antennas are placed surrounding the target, with each antenna sequentially transmitting signals, and all received signals are directly uploaded for centralized processing. Compared to multi-antenna EM imaging, the key differences in this paper are: (1) the signals received by the antennas are not directly uploaded to the CPU but instead, the sensing results are fused to reduce communication overhead; (2) not every antenna transmits signals sequentially, but rather, beamforming is applied at the transmitter to reduce the mutual coherence of the pilot patterns, thereby improving sensing accuracy.